Wildhorse Micro Brewers generated revenues of $13,577,000 with a 60 percent capital intensity ratio during the year ended September 30, 2017. Its net income was $814,620. With the introduction of a half dozen new specialty beers, management expects to grow sales by 15 percent next year. Assume that all costs vary directly with sales and that the firm maintains a dividend payout ratio of 70 percent.

What will be the EFN needed by this firm?

EFN $enter the external financing needed in dollars


If the company wants to raise no more than $750,000 externally and is not averse to adjusting its dividend payout policy, what will be the new dividend payout ratio? (Round answer to 2 decimal places, e.g. 5.27%.)

Dividend payout ratio enter the dividend payout ratio in percentages rounded to 2 decimal places %
Answer #1

EFN = Increase in total assets - Internal funds generated through retained earnings

= Capital intensity ratio * S0 * growth rate - Profit margin * Retention ratio * S0 *(1+g)

= 0.60 * 13577000 * 15% - 814620 /13577000 * (100%-70%) * 13577000 * (1+15%)

= 1221930 - 281043.90

= 940886.10

If the maximum debt to be raised is $750000 then the extra funding required has to be met through retained earnings. This is only possible if the company is able to reduce its dividend

Reduction in dividend = 940886.10 - 750000 = 190886.10

Net profit in next year = 814620 * (1+15%) = 936813

Dividend in next year before any adjustments = 936813 * 70% = 655769.10

Dividend after adjustment = 655769.10 - 190886.10 = 464883

Dividend payout ratio = 464883 / 936813

= 49.62%